# Observing Hot Holographic Quark Star With Gravitational Waves

### v1

@@ Abstract
We extract the equation of state for hot quark matter from a holographic $2+1$ flavor QCD model, which could form the core of a stable compact star. By adding a thin hadron shell, a new type of hybrid star is constructed. With the temperature serving as a parameter, the EoS varies and we obtain stable stars with the maximum mass of around 23 to 30 solar masses, and the compactness around $0.1$. The I-Love-Q-C relations are further discussed, and compared with the neutron star cases. These compact stars are candidates for black hole mimickers, which could be observed by gravitational waves and distinguished by properties like nonzero tidal Love number and electromagnetic signals.
@@ 1 Introduction
Inside the core of NS, the high pressure may cause the quark to be exposed, and many studies on quark stars (QS) have been performed. In parallel to the D4/D8 brane setup, there are studies on quark EoS based on D3/D7 model Hoyos:2016zke ; Hoyos:2016cob ; Annala:2017tqz , which leads to the conclusion that the quark matter cannot exist in neutron stars. Besides, the EoS itself is "softer" than the NS EoS, making the corresponding stars less compact than a normal NS, which is against the instinct. In Zhang:2022uin , within the EMD system and Karch-Katz-Son-Stephanov (KKSS) framework, parameters from model in Cai:2022omk are utilized to study the EoS for cold dense matter inside NS. It was also found that quark matter in NS is unstable. As such, for the current cold quark matter core models, the star is usually unstable, except for some studies on the strange quark matter stars Kurkela:2009gj .
@@
A way out is to consider another possibility. Rather than cold quark matter under high pressure, there could also be hot quark matter with lower pressure. In Aoki:2006we ; Bazavov:2009zn , studies on hot quark matter are performed, with the state equation and QCD phase transition given by lattice QCD method. In Zollner:2024iza , warm EoS is obtained from the EMD model. In Cai:2022omk , a holographic model for $2+1$ flavor QCD compatible with lattice QCD data is proposed (and in Zhao:2023gur the 2-flavor case), with finite chemical potential and finite temperature. In this paper, we extract the EoS for quark-gluon plasma phase from this 2 + 1 flavor model, which could form a quark core of a compact star. For the outer layer, it may cool down to hadron phase, where we could use a (hot) neutron-like EoS. We then discuss the mass-radius and I-Love-Q-C relations Yagi:2013awa ; Yagi:2013bca ; Yagi2014WhyIE of this two-layer model, following the method in Zhang:2020pfh . It is possible to interpret some GW events as this kind of hot QS, which on the contrary will constrain the parameters in this model.
@@
Since now we have two variables, the chemical potential $\mu$ and the temperature $T$, for simplicity, we choose five different constrains between the two, to illustrate the wide parameter space. We use fixed ratios of $\mu/T$, fixed $\mu$ and fixed $T$, around the critical end point (CEP). The maximum masses of the resultant stars lie between 23 to 30 $M_{\odot}$ (solar mass), with the compactness around $0.1$. These belong to a totally new kind of compact stars, very different from NS, mainly resulting from their lower central pressure and higher temperature. They may be formed through supernova-like process from massive stars more than 30 $M_{\odot}$, but the high temperature prevent them from collapsing into black holes (BH). The formation mechanism is beyond the range of the current paper, and will be left for future work.
@@
The potential observation of this kind of hot QS is hopeful. Judging from their masses, they could be detected as BH mimickers through GW observatories. But there is a significant difference from BH: their low compactness will lead to high TLN and different wave forms, rather than zero TLN for BH. Thus, it will be easy to distinguish hot QS. In this sense, it is meaningful to recheck the TLN in GW events with one companion less than 30 $M_{\odot}$, such as GW190413_052954, GW190630_185205, GW170814, GW190828_063405, GW200209_085452 PhysRevX.9.031040 ; PhysRevX.13.041039 and pay special attention to future events within this range. Furthermore, the high temperate quark matter in such hot QS will have characteristic EM effects (though its significance depends a lot on the distance from the earth), making it also very different from possible dark stars which are supposed to have hardly any EM signal.
@@ 2 Holographic QCD Model
The no-hair theorem is violated in asymptotic AdS spacetime Torii:2001pg , and the emergence of scalar hair contributes additional terms to the first law of BH thermodynamics Lu:2014maa . Research in Li:2020spf investigates the thermodynamics of AdS BH under Einstein-scalar gravity, defining thermodynamic quantities and mass through holographic renormalization. In Cai:2022omk scenarios with nonzero chemical potential are introduced, compatible with the lattice model with zero chemical potential. Our calculations directly employ the model in Cai:2022omk . Unlike previous studies focusing on cold quark matter, in this model we consider quark-gluon plasmas around the CEP, with the temperature reaching as high as 100 MeV ($1MeV\approx 1.16\times 10^{10}K$).
@@
$S=\frac{1}{2\kappa_{N}^{2}}\int d^{5}x\sqrt{-g}\left[R-\frac{1}{2}\bigtriangledown_{\mu}\phi\bigtriangledown^{\mu}\phi-\frac{Z(\phi)}{4}F_{\mu\nu}F^{\mu\nu}-V(\phi)\right],$ (1)
@@
with $b=-0.27341$, as a result of holographic renormalization. We can take the data on the horizon as an initial condition, solve the differential equation system with the NDSolve numerical method, then apply the numerical solution to fit the parameters.
@@
In QCD the common units of energy density and pressure are both $[MeV^{4}]$, when performing subsequent calculations, they need to be converted to astrophysical units: $r_{\odot}=G_{N}M_{\odot}/c^{2},\epsilon_{\odot}=M_{\odot}/r_{\odot}^{3}$, and $p_{\odot}=c^{2}\epsilon_{\odot}$, and we have
$MeV^{4}$ $=935340kg/m^{3}=1.51453*10^{-15}\epsilon_{\odot},$
$MeV^{4}$ $=8.40643*10^{22}kg/(s^{2}m)=1.51453*10^{-15}p_{\odot}.$ (5)
We can then extract the EoS: comparing to the cold quark models which usually use zero temperature, now $T$ serves as an extra parameter, together with the chemical potential $\mu$. We choose to fix the ratio between $T$ and $\mu$ to grab some general information of the parameter space, rather than discussing their detailed relation with the radius. Since the CEP obtained in Cai:2022omk is at $T=105MeV$ and $\mu=555MeV$, we pick five typical data, namely $\mu/T=4.5$, $\mu/T=5.28$, $\mu/T=6$, $T=105MeV$ and $\mu=555MeV$, to extract the corresponding EoS. Their energy density-pressure data graphs and corresponding fitting curves are drawn respectively in Fig.1.
@@
There are some nuances we cannot ignore. See Fig.1, for the quark phase, the pressure cannot drop to 0, but around $10^{-7}p_{\odot}$ (or about $10^{7}[MeV^{4}]$), where the phase transition lies. For the case of $T=105MeV$, $\mu=555MeV$, $\mu/T=5.28$ and $\mu/T=6$, they have obvious phase transition points, and the first three cases pass through CEP in Fig.2, so the CEP is directly used as the lower limit of the data ($p=3.97×10^{-8}p_{\odot}=2.62127\times 10^{7}MeV^{4}$). For $\mu/T=6$, it can be seen from Fig.3 that at 605MeV, the data point is on the left of the phase transition point, and the data point of 607MeV is on the right, so that the lower limit of the data can be read off ($p=3.03394×10^{-8}p_{\odot}=2.00322\times 10^{7}MeV^{4}$). Although $\mu/T=4.5$ does not pass through the first-order phase transition line, we find a distinct high order phase transition point in the pressure-energy data, so we take the data before that point ($p=6.07854×10^{-8}p_{\odot}=4.01348\times 10^{7}MeV^{4}$).
@@
Their fitting results ($p_{1},\epsilon_{1}$ corresponds to $\mu/T=4.5$, $p_{2},\epsilon_{2}$ corresponds to $\mu/T=5.28$, $p_{3},\epsilon_{3}$ corresponds to $\mu/T=6$, $p_{4},\epsilon_{4}$ corresponds to $T=105MeV$, $p_{5},\epsilon_{5}$ corresponds to $\mu=555MeV$ ) are listed respectively:
$\quad\quad\quad\epsilon_{1}$ $=0.0592293p_{1}^{0.651263}+1.43266\times 10^{13}p_{1}^{3.42233},$
$p_{1}$ $\in[6.07854\times 10^{-8},2.42143\times 10^{-6}],\epsilon_{1}\in[8.65502\times 10^{-7},0.000013846],$ (6)
$\epsilon_{2}$ $=0.027042p_{2}^{0.593438}+1.80685\times 10^{10}p_{2}^{2.87626},$
$p_{2}$ $\in[3.97089\times 10^{-8},2.5162\times 10^{-6}],\epsilon_{2}\in[7.09555\times 10^{-7},0.0000142],$ (7)
$\quad\quad\quad\epsilon_{3}$ $=0.000833022p_{3}^{0.370279}+8.72125p_{3}^{1.08721},$
$p_{3}$ $\in[3.03394\times 10^{-8},9.51759\times 10^{-6}],\epsilon_{3}\in[1.36385\times 10^{-6},0.000019245],$ (8)
$\>\>\quad\quad\quad\epsilon_{4}$ $=0.017254p_{4}^{0.557614}+3330.51p_{4}^{1.68025},$
$p_{4}$ $\in[3.97113\times 10^{-8},7.12588\times 10^{-6}],\epsilon_{4}\in[7.98295\times 10^{-7},0.0000305948],$ (9)
$\quad\quad\quad\epsilon_{5}$ $=0.0235768p_{5}^{0.590736}+23025.1p_{5}^{1.77162},$
$p_{5}$ $\in[3.97265\times 10^{-8},3.60026\times 10^{-6}],\epsilon_{5}\in[6.111821\times 10^{-7},0.00001945].$ (10)
@@ 3 Hot Quark Star
Although the minimum quark-gluon plasma phase is nonzero, the approach to zero pressure is essential to form a star, because the zero pressure represents the star surface. The most naive way to find the EoS to zero pressure is to simply use the fitting EoS, as extrapolations of the original data, which are valid only larger than the phase transition values. This means a smooth phase transition from the quark core to the outer hadron layer, not realistic enough, but can serve as a reference. In Fig.4, we present the EoS and the I-Love-Q relations for this kind of "simple" QS models (this terminology are used afterwards). Additionally, the M-R relation and the relations related to the compactness $C$ are shown, applying the calculation procedures in PhysRevD.88.023009 . We obtain that, the maximum masses are about 23 to 30 $M_{\odot}$, and the compactness is about $0.1$.
More rigorously, since there exists a phase transition from quark-gluon plasma core to the hadronic outer layer as the star cools, we should connect the holographic EoS derived above with a hot neutron-like model to calculate the mass-radius relation and the I-Love-Q-C relations. For the sake of simplicity in illustrating the idea, we adopt a single-component polytropic EoS to model the outer layer
$\epsilon_{ns}=\kappa_{ns}\,p^{\gamma_{ns}}.$ (11)
@@
$\epsilon_{ns}=n\,\epsilon_{qs},$ (12)
@@
In addition to the discontinuous EoS with $n=0.5$ in Fig.8, it already demonstrates that the I-Love-Q trio are broken even in the continuous model in Fig.4. Actually, the results from Wu:2023aaz indicate that the relations for extreme continuous EoS are also broken, in addition to the cases verified in many previous papers PhysRevD.88.023007 ; Haskell:2013vha ; Silva2017ILoveQTT ; PhysRevD.103.023022 ; Li:2023owg ; Atta:2024ckt ; Kumar:2023ojk ; Yagi2014WhyIE ; Pani:2015tga . Regarding the I-C, Love-C and Q-C relations, the universality is clearly lost. However, as pointed out in PhysRevD.88.023009 , the loss of universality in compactness relations for realistic NS also occurs, which lends support to the rationality of our findings.
@@
By comparing the panels for M-R relations in these figures, we observe that the hybrid stars become unstable for large phase transitions, as shown in Fig.8. Interestingly, the maximum M-R values for each model appear to match those of the "simple” quark case, though deviations become more pronounced when $\mu/T=4.5$. To investigate the cause, we calculate the maximum mass and radius at the phase transition point for each model (the quark core) in Table.1. The M-R pairs for the cores are listed in the second column, while the total star M-R pairs for "simple” QS are provided in the last column. It is evident that all models exhibit heavy cores (with the exception of $\mu/T=4.5$ case), and the M-R relations are significantly influenced by the core properties.
@@
The panels of I-Love-Q-C relations show that the universal relations are observed when applying connected EoS, with the exception of $\mu/T=4.5$. In PhysRevD.88.023009 , a possible explanation for the existence of I-Love-Q relations is suggested, namely that the EoS of realistic NS models approach each other in the outer layers. Since we connect the same polytropic models to various quark models, it is plausible that the deviations in the I-Love-Q-C relations for "simple” QS are reduced, even though these stars possess relatively light shells.
@@
In conclusion, the mass distribution of both "simple” QS and hybrid stars, all consisting of quark cores and hadron outer layers, allows these objects to mimic stellar-mass BH. The obedience of I-Love-Q-C relations for hybrid models presented in this paper extends the validity of these universal relations. On the front of observation, the I-Love-Q-C relations offer a potential means to distinguish between BH and QS because of the zero TLN for BH in GR. Furthermore, the broken relations about compactness help identify the holographic quark models. Since the I-Love-Q-C relations are valid for most of the hybrid models studied in this paper, these stars could potentially be distinguished from the "simple” quark cases. Additionally, the observation on the hybrid stars could place constraints on the phase transition from quark matter to hadronic matter when applying such models.
@@
Figure 4: The EoS, M-R relation and I-Love-Q-C relations for "simple” quark star with different models in eq.(6-10). For reference, the black dotted line representing a polytropic NS $\epsilon=0.09p^{0.5}$ is plotted in the panels of I-Love-Q-C relations.

### v2

# Predictions on observing hot holographic quark star with gravitational waves

@@ Abstract
We extract the equation of state of hot quark matter from a holographic $2+1$ flavor QCD model, which could form the core of a stable compact star. By adding a thin hadron shell, a new type of hybrid star is constructed. With the temperature serving as a parameter, the EoS varies and we obtain stable stars with mass ranging from about 5 to 30 solar masses, and the maximum compactness around $0.2$. The I-Love-Q-C relations are further discussed, and compared with the neutron star cases. These compact stars are candidates for black hole mimickers, which could be observed by gravitational waves and distinguished by properties like nonzero tidal Love number and electromagnetic signals.
journal: Eur. Phys. J. C
@@ 1 Introduction
Inside the core of NS, the high pressure may cause the quark to be exposed, and many studies on quark stars (QS) have been performed. In parallel to the D4/D8 brane setup, there are studies on quark EoS based on D3/D7 model Hoyos:2016zke ; Hoyos:2016cob ; Annala:2017tqz ; BitaghsirFadafan:2019ofb , which leads to the conclusion that the quark matter cannot exist in neutron stars. Besides, the EoS itself is "softer" than the NS EoS, making the corresponding stars less compact than a normal NS, which is against the instinct. In Zhang:2022uin , within the EMD system and Karch-Katz-Son-Stephanov (KKSS) framework, parameters from model in Cai:2022omk are utilized to study the EoS for cold dense matter inside NS. It was also found that quark matter in NS is unstable. There are indeed results on hybrid stars with stable cold quark cores with special setup BitaghsirFadafan:2020otb , or predicted by observations and the sound speed Annala:2019puf ; Han:2022rug ; Fan:2023spm , and also some studies on stable strange quark matter stars Kurkela:2009gj .
@@
A way out is to consider another possibility. Rather than cold quark matter under high pressure, there could also be hot quark matter with lower pressure. In Aoki:2006we ; Bazavov:2009zn , studies on hot quark matter are performed, with the state equation and QCD phase transition given by lattice QCD method. In Zollner:2024iza , warm EoS is obtained from the EMD model. In Cai:2022omk , a holographic model for $2+1$ flavor QCD compatible with lattice QCD data is proposed (and in Zhao:2023gur the 2-flavor case), with finite chemical potential and finite temperature. In this paper, we extract the EoS for quark-gluon plasma phase from this 2 + 1 flavor model, which could form a quark core of a compact star. For the crust, it may cool down to hadron phase, where we could use a (hot) neutron-like EoS. We then discuss the mass-radius and I-Love-Q-C relations Yagi:2013awa ; Yagi:2013bca ; Yagi2014WhyIE of this two-layer model, following the method in Zhang:2020pfh . It is possible to interpret some GW events as this kind of hot QS, which on the contrary will constrain the parameters in this model.
@@
Since now we have two variables, the chemical potential $\mu$ and the temperature $T$, for simplicity, we choose five different constrains between the two, to illustrate the wide parameter space. We use fixed ratios of $\mu/T$, fixed $\mu$ and fixed $T$, around the critical end point (CEP). The maximum masses of the resultant stars lie between 23 to 30 $M_{\odot}$ (solar mass), with the compactness around $0.2$. The minimum masses are about 5 $M_{\odot}$. These belong to a totally new kind of compact stars, very different from NS, mainly resulting from their lower central pressure and higher temperature.
The potential observation of this kind of hot QS is hopeful. Judging from their masses, they could be detected as black hole (BH) mimickers Lemos:2008cv ; Cardoso:2019rvt through GW observatories. BH mimickers are exotic compact objects that imitate many observable properties of BH such as extreme compactness, strong gravitational fields, and sometimes similar GW or electromagnetic signatures, but do not possess event horizons. They could be boson stars Liebling:2012fv , gravastars Mazur:2001fv , fuzzballsMathur:2005zp or QS, etc, since their existence cannot be easily excluded from observational data in favor of BH Abramowicz:2002vt .
Hot QS has a significant difference from BH: the low compactness will lead to high TLN and different wave forms, rather than zero TLN for BH. Thus, it will be not difficult to distinguish hot QS if one could gain access to the TLN information, which GW interferometers like LIGO/Virgo could provide LIGOScientific:2017vwq . In binary systems of compact stars, the GW waveform encodes the tidal deformability of the components. Instead of measuring the individual TLN of the two stars, a combined tidal deformability parameter defined as
@@
$\tilde{\Lambda}=\frac{16}{13}\frac{(M_{1}+12M_{2})M_{1}^{4}\Lambda_{1}+(M_{2}+12M_{1})M_{2}^{4}\Lambda_{2}}{(M_{1}+M_{2})^{5}}$ is accessible, where $M_{1}$ and $M_{2}$ are the masses of the two stars, and $\Lambda_{1}$ and $\Lambda_{2}$ are their respective TLN Flanagan_2008 ; Hinderer_2010 . The tidal deformability of a hot QS–NS binary differs markedly from that of a BH–NS binary with the same NS companion. For example, for a binary consisting of a 20 $M_{\odot}$ BH and a 2 $M_{\odot}$ NS (with $\Lambda\approx 20$), the combined deformability is $\tilde{\Lambda}=0.02$, which is likely undetectable. In contrast, replacing the BH with a 20 $M_{\odot}$ hot QS (with $\Lambda\approx 1000$ as an example) yields $\tilde{\Lambda}=1681$. Such a large difference in tidal response would make it possible to distinguish QS from BH. In this sense, it is meaningful to recheck the TLN in GW events with one companion less than 30 $M_{\odot}$, such as GW190413_052954, GW190630_185205, GW170814, GW190828_063405, GW200209_085452 PhysRevX.9.031040 ; PhysRevX.13.041039 and pay special attention to future events within this range. Furthermore, the high temperate quark matter in such hot QS will have characteristic EM effects (though its significance depends a lot on the distance from the earth), making it also very different from possible dark stars which are supposed to have hardly any EM signal.
@@ 2 Holographic QCD Model
The no-hair theorem is violated in asymptotic AdS spacetime Torii:2001pg , and the emergence of scalar hair contributes additional terms to the first law of BH thermodynamics Lu:2014maa . InGubser:2008yx , a five-dimensional gravity model with a scalar field was considered, the potential of the scalar field is designed to reproduce the equation of state of QCD. Based on this, DeWolfe:2010he ; DeWolfe:2011ts simulates the phase transitions and critical phenomena of QCD at finite temperature and density. To construct the equation of state for neutron stars, we require the explicit forms of energy and pressure. Therefore, we chose another holographic QCD model. Research in Li:2020spf investigates the thermodynamics of AdS BH under Einstein-scalar gravity, defining thermodynamic quantities and mass through holographic renormalization. In Cai:2022omk scenarios with nonzero chemical potential are introduced, compatible with the lattice model with zero chemical potential. Our calculations directly employ the model in Cai:2022omk . Unlike previous studies focusing on cold quark matter, in this model we consider quark-gluon plasmas around the CEP, with the temperature reaching as high as 100 MeV ($1\rm{MeV}\approx 1.16\times 10^{10}\rm{K}$).
@@
$\begin{split}S=&\frac{1}{2\kappa_{N}^{2}}\int d^{5}x\sqrt{-g}\\ &\left[R-\frac{1}{2}\bigtriangledown_{\mu}\phi\bigtriangledown^{\mu}\phi-\frac{Z(\phi)}{4}F_{\mu\nu}F^{\mu\nu}-V(\phi)\right],\end{split}$ (1)
@@
Comparing to the original form in DeWolfe:2010he , the form of $Z(\phi)$ exhibits a spurious spike at low $\phi$. This is due to the presence of the last term in 2, which is numerically zero. In Jokela:2024xgz , it is suggested that this term could be safely removed to simplify the analysis. Here we stay with this form of $Z(\phi)$ in Cai:2022omk , since it not only is obtained by comparing with lattice data consistently, but also can explicitly provides the energy density and pressure. Besides, although this spike may effect the results at high densities and low temperatures (corresponding to low $\phi$), in this paper we are dealing with the opposite region.
@@
with b=0.27341b=-0.27341, as a result of holographic renormalization. We can take the data on the horizon as an initial condition, solve the differential equation system with the NDSolve numerical method, then apply the numerical solution to fit the parameters.
@@
In QCD the common units of energy density and pressure are both $\rm{MeV}^{4}]$, when performing subsequent calculations, they need to be converted to astrophysical units: $r_{\odot}=G_{N}M_{\odot}/c^{2},\epsilon_{\odot}=M_{\odot}/r_{\odot}^{3}$, and $p_{\odot}=c^{2}\epsilon_{\odot}$, and we have
$\rm{MeV}^{4}$ $=935340\rm{kg}/\rm{m}^{3}=1.51453\times 10^{-15}\epsilon_{\odot},$
$\rm{MeV}^{4}$ $=8.40643\times 10^{22}\rm{kg}/(\rm{s}^{2}\rm{m})=1.51453\times 10^{-15}\textit{p}_{\odot}.$ (5)
We can then extract the EoS: comparing to the cold quark models which usually use zero temperature, now $T$ serves as an extra parameter, together with the chemical potential $\mu$. Since the specific relationship for $\mu$ and T as functions of the position in the star is complicated and unknown, we choose to fix the ratio between $T$ and $\mu$ to grab some general information of the parameter space, rather than discussing their detailed relation with the radius. Since the CEP obtained in Cai:2022omk is at $T=105\rm{MeV}$ and $\mu=555\rm{MeV}$, we pick five typical data, namely $\mu/T=4.5$, $\mu/T=5.28$, $\mu/T=6$, $T=105\rm{MeV}$ and $\mu=555\rm{MeV}$, to extract the corresponding EoS. For the data where T is fixed, we chose $\mu\in[555\rm{MeV},1440\rm{MeV}]$, for the data where $\mu$ is fixed, we chose $T\in[105\rm{MeV},180\rm{MeV}]$. Their energy density-pressure data graphs and corresponding fitting curves are drawn respectively in Fig.1.
@@
There are some nuances we cannot ignore. See Fig.1, for the quark phase, the pressure cannot drop to 0, but around $10^{-7}p_{\odot}$ (or about $10^{7}[\rm{MeV}^{4}]$), where the phase transition (quark-gluon plasma phase transition to hadrons) lies. For the case of $T=105\rm{MeV}$, $\mu=555\rm{MeV}$, $\mu/T=5.28$ and $\mu/T=6$, they have obvious phase transition points, and the first three cases pass through CEP in Fig.2, so the CEP is directly used as the lower limit of the data ($p=3.97\times 10^{-8}p_{\odot}=2.62127\times 10^{7}\rm{MeV}^{4}$). For $\mu/T=6$, it can be seen from Fig.3 that at 605MeV, the data point is on the left of the phase transition point, and the data point of 607MeV is on the right, so that the lower limit of the data can be read off ($p=3.03394\times 10^{-8}p_{\odot}=2.00322\times 10^{7}\rm{MeV}^{4}$). Although $\mu/T=4.5$ does not pass through the first-order phase transition line, we find a distinct high order phase transition point in the pressure-energy data, so we take the data before that point ($p=6.07854\times 10^{-8}p_{\odot}=4.01348\times 10^{7}\rm{MeV}^{4}$).
@@
Their fitting results ($p_{1},\epsilon_{1}$ corresponds to $\mu/T=4.5$, $p_{2},\epsilon_{2}$ corresponds to $\mu/T=5.28$, $p_{3},\epsilon_{3}$ corresponds to $\mu/T=6$, $p_{4},\epsilon_{4}$ corresponds to $T=105\rm{MeV}$, $p_{5},\epsilon_{5}$ corresponds to $\mu=555\rm{MeV}$ ) are listed, respectively:
$\begin{split}\epsilon_{1}&=5.92293\times 10^{-2}\,p_{1}^{0.651263}+1.43266\times 10^{13}\,p_{1}^{3.42233},\\ p_{1}&\in[6.07854\times 10^{-8},\;2.42143\times 10^{-6}],\\ \epsilon_{1}&\in[8.65502\times 10^{-7},\;1.3846\times 10^{-5}],\end{split}$ (6)
$\begin{split}\epsilon_{2}&=2.7042\times 10^{-2}\,p_{2}^{0.593438}+1.80685\times 10^{10}\,p_{2}^{2.87626},\\ p_{2}&\in[3.97089\times 10^{-8},\;2.5162\times 10^{-6}],\\ \epsilon_{2}&\in[7.09555\times 10^{-7},\;1.42\times 10^{-5}],\end{split}$ (7)
$\begin{split}\epsilon_{3}&=8.33022\times 10^{-4}\,p_{3}^{0.370279}+8.72125\,p_{3}^{1.08721},\\ p_{3}&\in[3.03394\times 10^{-8},\;9.51759\times 10^{-6}],\\ \epsilon_{3}&\in[1.36385\times 10^{-6},\;1.9245\times 10^{-5}],\end{split}$ (8)
$\begin{split}\epsilon_{4}&=1.7254\times 10^{-2}\,p_{4}^{0.557614}+3.33051\times 10^{3}\,p_{4}^{1.68025},\\ p_{4}&\in[3.97113\times 10^{-8},\;7.12588\times 10^{-6}],\\ \epsilon_{4}&\in[7.98295\times 10^{-7},\;3.05948\times 10^{-5}],\end{split}$ (9)
$\begin{split}\epsilon_{5}&=2.35768\times 10^{-2}\,p_{5}^{0.590736}+2.30251\times 10^{4}\,p_{5}^{1.77162},\\ p_{5}&\in[3.97265\times 10^{-8},\;3.60026\times 10^{-6}],\\ \epsilon_{5}&\in[6.111821\times 10^{-7},\;1.945\times 10^{-5}].\end{split}$ (10)
@@
Again we must emphasize that, none of the above 5 cases shows the real EoS for the QS, but they constrain the range where the real EoS could lie in.
@@ 3 Hot Quark Star
Although the minimum quark-gluon plasma phase is nonzero, the approach to zero pressure is essential to form a star, because the zero pressure represents the star surface. The most naive way to find the EoS to zero pressure is to simply use the fitting EoS, as extrapolations of the original data, which are valid only larger than the phase transition values. This means a smooth phase transition from the quark core to the outer hadron layer, not realistic enough, but can serve as a reference. In Fig.4, we present the EoS and the I-Love-Q relations for this kind of "simple" QS models (this terminology are used afterward). Additionally, the M-R relation and the relations related to the compactness $C$ are shown, applying the calculation procedures in PhysRevD.88.023009 . We obtain that, the maximum masses are about 23 to 30 $M_{\odot}$, and the compactness is about $0.1$.
More rigorously, since there exists a phase transition from quark-gluon plasma core to the hadronic crust as the star cools, we should connect the holographic EoS derived above with a hot (the temperature is expected to drop toward the surface) neutron-like model to calculate the mass-radius relation and the I-Love-Q-C relations. For the sake of simplicity in illustrating the idea, we adopt a single-component polytropic EoS to model the outer layer
$\qquad\qquad\qquad\qquad\epsilon_{ns}=\kappa_{ns}\,p^{\gamma_{ns}}.$ (11)
@@
$\qquad\qquad\qquad\qquad\epsilon_{ns}=n\,\epsilon_{qs},$ (12)
@@
Notice that all the curves (the M-R relations etc.) in Fig.5-8 (except the left-top panels for EoS) are illustrated with the pressure ranges given in (6) to (10) , which means the lowest star central pressures lie on the phase transition points from quark-gluon to hadrons. Below those chosen lowest pressure values, the stars will be no longer QS but pure hadronic, without any quark-gluon state, and in our case the corresponding EoS for pure hadrons are just single polytropic, whose behaviors are well-known. All the stars represented in the curves have quark cores, while stars made by pure hadrons are not drawn, so no apparent phase transitions are shown, in contrast to those in the left-top panel EoS curves.
@@
In addition to the discontinuous EoS with $n=0.5$ in Fig.8, it already demonstrates that the I-Love-Q trio are broken even in the continuous model in Fig.4. Actually, the results from Wu:2023aaz indicate that the relations for extreme continuous EoS are also broken, in addition to the cases verified in many previous papers Yagi2014WhyIE ; Haskell:2013vha ; Silva2017ILoveQTT ; PhysRevD.103.023022 ; Li:2023owg ; Atta:2024ckt ; Kumar:2023ojk ; Pani:2015tga . Regarding the I-C, Love-C and Q-C relations, the universality is clearly lost. However, as pointed out in PhysRevD.88.023009 , the loss of universality in compactness relations for realistic NS occurs. Also in Hoyos:2021uff , the authors analytically obtain the quasi-universal relations of a holographic quark eos, which lends support to the rationality of our findings.
@@
By comparing the panels for M-R relations in these figures, we observe that the hybrid stars become unstable for large phase transitions, as shown in Fig.8. Interestingly, the maximum M-R values for each model appear to match those of the "simple" quark case, though deviations become more pronounced when $\mu/T=4.5$. To investigate the cause, we calculate the maximum mass and radius at the phase transition point for each model (the quark core) in Table.1. The M-R pairs for the cores are listed in the second column, while the total star M-R pairs for "simple" QS are provided in the last column. It is evident that all models exhibit heavy cores (with the exception of $\mu/T=4.5$ case), and the M-R relations are significantly influenced by the core properties.
@@
The panels of I-Love-Q-C relations show that the universal relations are observed when applying connected EoS, with the exception of $\mu/T=4.5$. In PhysRevD.88.023009 , a possible explanation for the existence of I-Love-Q relations is suggested, namely that the EoS of realistic NS models approach each other in the outer layers. Since we connect the same polytropic models to various quark models, it is plausible that the deviations in the I-Love-Q-C relations for "simple" QS are reduced, even though these stars possess relatively light shells.
@@
In conclusion, the mass distribution of both "simple" QS and hybrid stars, all consisting of quark cores and hadron outer layers, allows these objects to mimic stellar-mass BH. The obedience of I-Love-Q-C relations for hybrid models presented in this paper extends the validity of these universal relations. On the front of observation, the I-Love-Q-C relations offer a potential means to distinguish between BH and QS because of the zero TLN for BH in GR. Furthermore, the broken relations about compactness help identify the holographic quark models. Since the I-Love-Q-C relations are valid for most of the hybrid models studied in this paper, these stars could potentially be distinguished from the "simple" quark cases. Additionally, the observation on the hybrid stars could place constraints on the phase transition from quark matter to hadronic matter when applying such models.
@@
Figure 4: The EoS, M-R relation and I-Love-Q-C relations for "simple" quark star with different models in eq.(6-10). For reference, the black dotted line representing a polytropic NS $\epsilon=0.09p^{0.5}$ is plotted in the panels of I-Love-Q-C relations.